# Integrating Dynamical Systems Modeling with Spatiotemporal scRNA-seq Data Analysis

### v1

@@ 3.1.3 Continuous Dynamics Modeling
Several methods leverage the continuity assumption in single-cell data: If the observed data fully capture the continuous dynamics of cellular evolution, a cell’s state at the next timestep should align with its neighbors. Several methods use such prior for defining the loss function to further refine the RNA velocity.
DeepVelo (Cui et al., 2024) parameterizes $\alpha_{g,c},\beta_{g,c},\gamma_{g,c}$ via neural networks to compute RNA velocity $v_{g,c}$. Its loss function incorporates temporal consistency: The state $\mathbf{s}_{c}(t+1)$ is approximated as a weighted sum of neighboring cell states $\mathbf{s}_{j}(t(j))$, where the transition probability $\mathbb{P}_{+}(c\rightarrow j)$ is defined as:
$\mathbb{P}_{+}(c\rightarrow j)=\begin{cases}\dfrac{1}{Z}&\text{if }\cos(\mathbf{s}_{j}-\mathbf{s}_{c},\mathbf{\hat{v}}_{c})>0\ \text{and}\ j\in\mathcal{N}(c),\\ 0&\text{otherwise}.\end{cases}$
where $Z$ is the normalization constant. The forward consistency loss enforces this assumption:
$\mathcal{L}_{+}=\sum_{c=1}^{n_{c}}\left\|\mathbf{s}_{c}(t)+\mathbf{\hat{v}}_{c}-\mathbf{s}_{c}(t+1)\right\|^{2},$
with $\mathbf{s}_{c}(t+1)=\sum_{j\in\mathcal{N}(c)}\mathbf{s}_{j}(t(j))\mathbb{P}_{+}(c\rightarrow j)$. A symmetric backward consistency loss $\mathcal{L}_{-}$ is similarly constructed. Additionally, a correlation loss ensures consistency with transcriptional dynamics:
$\mathcal{L}_{\mathrm{corr}}=-\left[\lambda_{u}\dfrac{\mathbf{\hat{v}}_{c}\cdot\mathbf{u}_{c}}{\|\mathbf{\hat{v}}_{c}\|\|\mathbf{u}_{c}\|}+\lambda_{s}\dfrac{\mathbf{\hat{v}}_{c}\cdot\mathbf{s}_{c}}{\|\mathbf{\hat{v}}_{c}\|\|\mathbf{s}_{c}\|}\right],$
@@
CellDancer (Li et al., 2024b) similarly parameterizes $\alpha_{g,c},\beta_{g,c},\gamma_{g,c}$ with neural networks, and estimates velocity as:
$\mathbf{\hat{v}}_{c}=(\Delta\mathbf{u}_{c},\Delta\mathbf{s}_{c}),\quad\text{where }\Delta\mathbf{u}_{c}=\mathbf{\alpha}_{c}-\mathbf{\beta}_{c}\odot\mathbf{u}_{c},\ \Delta\mathbf{s}_{c}=\mathbf{\beta}_{c}-\mathbf{\gamma}_{c}\odot\mathbf{s}_{c}.$
Its loss maximizes velocity alignment with neighbors
$\mathcal{L}=\sum_{c}\left[1-\max_{j\in\mathcal{N}(c)}\cos\left(\mathbf{\hat{v}}_{c},\mathbf{v}_{j}\right)\right],\quad\mathbf{v}_{j}=(\mathbf{u}_{j}-\mathbf{u}_{c},\mathbf{s}_{j}-\mathbf{s}_{c}).$
Vector Field Reconstruction based on RNA Velocity
After estimating the model parameters and obtaining the RNA velocity for each cell, an important downstream task is to predict the fate of cells based on the estimated velocity, e.g. the continuous differentiation trajectory of cells. To construct a continuous vector field, the RNA velocity of each cell could be treated as the value of a vector field at discrete points. Then the vector field is reconstructed by formulating a regression problem. Subsequently, quantities derived from vector field analysis, such as equilibrium points, streamlines, gradients, divergence, and curl, are used to study cell fate in continuous dynamics setup (Qiu et al., 2022a).
@@
Denote the spliced RNA counts and RNA velocities at various data points $\{\mathbf{x}_{i},\mathbf{v}_{i}\}_{i=1}^{n}$. The task is to find a continuous vector field function $f^{\star}$ that minimizes the regression loss $\mathcal{L}=\sum_{i}p_{i}\|\mathbf{v}_{i}-\mathbf{f}(\mathbf{x}_{i})\|^{2}$ where $p_{i}$ denotes the weights of each data point.
Dynamo (Qiu et al., 2022a) approximates the unknown vector-valued function in a sparse reproducing kernel Hilbert space (RKHS). For a vector-valued function $\mathbf{f}\in\mathcal{H}$ in RKHS, it can be represented as a sum of Gaussian kernels
$\mathbf{f}(\mathbf{x})=\sum_{i=1}^{m}\Gamma(\mathbf{x},\mathbf{\tilde{x}}_{i})\mathbf{c}_{i},\quad\Gamma(\mathbf{x},\mathbf{\tilde{x}})=\exp(-w\|\mathbf{x}-\mathbf{\tilde{x}}\|^{2}),$
where $\mathbf{\tilde{x}}_{i}$ are called control points. Additionally, the norm of $\mathbf{f}$ in $\mathcal{H}$ can be computed as $\|\mathbf{f}\|^{2}=\sum_{i,j=1}^{m}\mathbf{c}_{i}^{T}\Gamma(\mathbf{\tilde{x}}_{i},\mathbf{\tilde{x}}_{j})\mathbf{c}_{j}^{T}.$ The loss function for the vector field estimation problem includes a regularization term based on the norm of $\mathbf{f}$ such that $\mathcal{L}_{\lambda}=\sum_{i}p_{i}\|\mathbf{v}_{i}-\mathbf{f}(\mathbf{x}_{i})\|^{2}+\frac{\lambda}{2}\|\mathbf{f}\|^{2},$ where $\lambda$ is the regularization coefficient.
Another popular fitting strategy to reconstruct the continuous vector field is to use the neural network, where a VAE-based deep learning method was proposed in (Chen et al., 2022d).
@@
Based on the estimated continuous vector field, Dynamo (Qiu et al., 2022a) proposed several analyses to reveal the differential geometry of the RNA velocity. First, the Jacobian matrix is essential for analyzing the stability of equilibrium points in a dynamical system and for studying gene-gene interactions. In RKHS context, the Jacobian matrix can be analytically computed as
$\mathbf{J}=\frac{\partial\mathbf{f}(\mathbf{x})}{\partial\mathbf{x}}=-2w\sum_{i=1}^{m}\Gamma(\mathbf{x},\mathbf{\tilde{x}}_{i})\mathbf{c}_{i}(\mathbf{x}-\mathbf{\tilde{x}}_{i})^{T}$
where $\Gamma(\mathbf{x},\mathbf{\tilde{x}}_{i})$ is the Gaussian kernel. Since the $(i,j)$-th entry of the Jacobian matrix $J_{ij}$ represents the effect of unspliced RNA levels of gene $j$ on the RNA velocity of gene $i$, the Jacobian matrix can be used to analyze the strength of gene-gene interactions. By averaging the Jacobian matrix across all data points, an average Jacobian matrix $\langle J\rangle$ can be obtained. By sorting the elements in each row of $\langle J\rangle$, the top regulators for each effector can be identified. Conversely, by sorting the elements in each column of $\langle J\rangle$, the top effectors for each regulator can be identified. The Jacobian matrix can also be used to compute the effect of perturbations. If the system state changes by $\Delta\mathbf{x}$ at $\mathbf{x}$, the resulting change in the vector field is $\Delta\mathbf{f}=\mathbf{J}\Delta\mathbf{x}.$
@@
    $\nabla\cdot\mathbf{f}=\mathrm{Tr}(\mathbf{J}),$
@@
    $\mathbf{a}=\frac{\mathrm{d}\mathbf{v}}{\mathrm{d}t}=\mathbf{J}\mathbf{v}.$
@@
    $\mathbf{\kappa}=\frac{1}{\|\mathbf{v}\|}\frac{\mathrm{d}}{\mathrm{d}t}\frac{\mathbf{v}}{\|\mathbf{v}\|}=\frac{(\mathbf{v}^{T}\mathbf{v})\mathbf{J}\mathbf{v}-(\mathbf{v}^{T}\mathbf{J}\mathbf{v})\mathbf{v}}{\|\mathbf{v}\|^{4}}.$
@@
Based on the learned vector field, continuous cell trajectories could also be constructed in Dynamo (Qiu et al., 2022a) using the concept of most probable path (Zhou and Li, 2016). Given the SDE (2.2), the action along any path $\Psi$ is defined as
[…]
$L^{FW}\left(\psi,\dot{\psi}\right)$ $=\frac{1}{4}\left[\dot{\psi}(s)-b\left(\psi(s)\right)\right]^{t}D^{-1}\left(\psi(s)\right)\left[\dot{\psi}(s)-b\left(\psi(s)\right).\right]$
According to the Freidlin–Wentzell theorem (Zhou and Li, 2016), the path of least action is indeed the most probable path to make transitions between two attractors. In actual computation, Dynamo assumes the constant noise coefficient by taking $D=\frac{\sigma^{2}}{2}$ since only a continuous vector field is reconstructed.
@@
$R(\mathbf{x}_{s}\rightarrow\mathbf{x}_{t})\approx C\exp(-S^{\star}),$ (3.6)
@@ 3.2 Temporally Resolved Single-cell RNA-seq
Temporally resolved scRNA-seq provides us with a deeper understanding of the dynamics process in single cells. However, due to the destructive nature of scRNA-seq technology, we cannot track the trajectories of individual cells. Instead, we can only observe the changes in cellular distribution with time. Thus, reconstructing the trajectories of single cells from samples collected at discrete and sparse temporal points becomes crucial for understanding developmental processes and other dynamic biological processes and remains a challenging problem (Schiebinger et al., 2019; Sha et al., 2024; Peng et al., 2024; Lavenant et al., 2024; Jiang et al., 2022; Jiang and Wan, 2024; Bunne et al., 2024, 2023b; Tong et al., 2023; Zhang et al., 2021; Maddu et al., 2024; Eyring et al., 2024; Zhang et al., 2024a; Yeo et al., 2021). To overcome these challenges, many methods have been developed in recent years. From a dynamical perspective, these approaches can be broadly classified into two categories: those that model dynamics on discrete cell states and those that model dynamics in continuous spaces. In the following, we will introduce these methods separately from these two viewpoints.
@@ 3.2.1 Discrete Temporal Dynamics Modeling
Among the methods that model dynamics on discrete cell states, pioneering work includes Waddington OT (Schiebinger et al., 2019) and Moscot (Klein et al., 2025). These approaches employ static optimal transport as the main tools.
Static Optimal Transport
To formulate this problem, consider $\mathbf{X}\in\mathbb{R}^{N\times G}$ and $\mathbf{Y}\in\mathbb{R}^{M\times G}$ represent two unpaired datasets of $N$ and $M$ cells observed at different time points ($t_{1},t_{2}$) respectively in the $G$ dimensional gene expression space. Then one can define two marginal distributions $\nu_{0}\in\mathcal{C}_{N}$, $\nu_{1}\in\mathcal{C}_{M}$ at $t_{1}$ and $t_{2}$ respectively on the probability simplex $\mathcal{C}_{N}=\{\mathbf{a}\in\mathbb{R}^{N}|\sum_{i=1}^{N}a_{i}=1,\mathbf{a}\geq 0\}.$ The goal of optimal transport is to find the optimal coupling $\boldsymbol{\pi}\in\mathbb{R}_{+}^{N\times M}$ that transports a distribution to another while minimizing the cost associated with the transportation. The feasible transport plan is defined as $\mathbf{\boldsymbol{\pi}}(\nu_{0},\nu_{1})=\left\{\boldsymbol{\pi}\in\mathbb{R}^{N\times M}:\boldsymbol{\pi}\mathbf{1}_{M}=\nu_{0},\boldsymbol{\pi}^{T}\mathbf{1}_{N}=\nu_{1},\boldsymbol{\pi}\geq 0\right\}.$ So the static optimal transport problem is formally defined as:
$\min_{\boldsymbol{\pi}\in\boldsymbol{\pi}(\nu_{0},\nu_{1})}\langle\boldsymbol{\pi},\mathbf{c}\rangle=\sum_{i,j}c_{i,j}\pi_{i,j}.$
The cost matrix $\mathbf{c}\in\mathbb{R}_{+}^{n\times m}$ defines the transportation cost between each pair of points, where $c_{ij}:=c(\mathbf{x}_{i},\mathbf{y}_{j})$ quantifies the expense of transferring a unit mass from the source point $\mathbf{x}_{i}$ to the target point $\mathbf{y}_{j}$. By solving the static optimal transport problem, one can determine the transport matrix that couples the two distributions. This static optimal transport can be effectively addressed using the Python Optimal Transport (POT) library (Flamary et al., 2021).
Entropic Optimal Transport
Additionally, to enhance the efficiency of solving the optimal transport problem, a regularized optimal transport approach is often introduced. The discrete entropy of a coupling matrix is defined as
@@
$\min_{\boldsymbol{\pi}\in\boldsymbol{\pi}(\mathbf{\nu_{0}},\mathbf{\nu_{1}})}\langle\boldsymbol{\pi},\mathbf{c}\rangle-\varepsilon H(\boldsymbol{\pi})$ (3.7)
Since the objective is an $\varepsilon$-strongly convex function, problem (3.7) has a unique optimal solution. Using the KKT conditions, the solution to (3.7) is unique and has the form (Peyré et al., 2019)
@@
for two unknown variables $(\mathbf{a},\mathbf{b})\in\mathbb{R}_{+}^{N}\times\mathbb{R}_{+}^{M}$, where $K_{i,j}=e^{-\frac{c_{i,j}}{\varepsilon}}$.
Sinkhorn algorithm From above, the optimal solution of problem (3.7) can be expressed in matrix form as $\boldsymbol{\pi}=\operatorname{diag}(\mathbf{a})\mathbf{K}\operatorname{diag}(\mathbf{b})$. Then it is necessary to satisfy the constraints $\boldsymbol{\pi}(\nu_{0},\nu_{1})$, i.e.,
$\operatorname{diag}(\mathbf{a})\mathbf{K}\operatorname{diag}(\mathbf{b})\mathbf{1}_{M}=\nu_{0},\quad\operatorname{diag}(\mathbf{a})\mathbf{K}^{T}\operatorname{diag}(\mathbf{b})\mathbf{1}_{N}=\nu_{1}.$
Note that $\operatorname{diag}(\mathbf{b})\mathbf{1}_{M}$ is $\mathbf{b}$ and the multiplication of $\operatorname{diag}(\mathbf{a})$ times $\mathbf{K}\mathbf{b}$ is
$\mathbf{a}\odot(\mathbf{K}\mathbf{b})=\nu_{0},\mathbf{a}\odot\left(\mathbf{K}^{\mathrm{T}}\mathbf{b}\right)=\nu_{1},$
where $\odot$ represents the entry-wise multiplication. An intuitive way is to solve them iteratively, and these two updates yield Sinkhorn’s algorithm,
$\mathbf{a}^{(\ell+1)}\stackrel{{\scriptstyle\text{ def. }}}{{=}}\frac{\nu_{0}}{\mathbf{K}\mathbf{b}^{(\ell)}},\quad\quad\mathbf{b}^{(\ell+1)}\stackrel{{\scriptstyle\text{ def. }}}{{=}}\frac{\nu_{1}}{\mathbf{K}^{{T}}\mathbf{a}^{(\ell+1)}}.$

### v2

@@ 3.1.3 Continuous Dynamics Modeling
Several methods leverage the continuity assumption in single-cell data: If the observed data fully capture the continuous dynamics of cellular evolution, a cell’s state at the next timestep should align with its neighbors. Several methods use such prior information for defining the loss function to further refine the RNA velocity.
DeepVelo Cui et al. (2024) parameterizes $\alpha_{g,c},\beta_{g,c},\gamma_{g,c}$ via neural networks to compute RNA velocity $v_{g,c}$. Its loss function incorporates temporal consistency: The state $\boldsymbol{s}_{c}(t+1)$ is approximated as a weighted sum of neighboring cell states $\boldsymbol{s}_{j}(t(j))$, where the transition probability $\mathbb{P}_{+}(c\rightarrow j)$ is defined as
$\mathbb{P}_{+}(c\rightarrow j)=\begin{cases}\dfrac{1}{Z}&\text{if }\cos(\boldsymbol{s}_{j}-\boldsymbol{s}_{c},\boldsymbol{\hat{v}}_{c})>0\ \text{and}\ j\in\mathcal{N}(c),\\ 0&\text{otherwise}.\end{cases}$
where $Z$ is the normalization constant. The forward consistency loss enforces this assumption as follows:
$\mathcal{L}_{+}=\sum_{c=1}^{n_{c}}\left\|\boldsymbol{s}_{c}(t)+\boldsymbol{\hat{v}}_{c}-\boldsymbol{s}_{c}(t+1)\right\|^{2},$
with $\boldsymbol{s}_{c}(t+1)=\sum_{j\in\mathcal{N}(c)}\boldsymbol{s}_{j}(t(j))\mathbb{P}_{+}(c\rightarrow j)$. A symmetric backward consistency loss $\mathcal{L}_{-}$ is similarly constructed. Additionally, a correlation loss ensures consistency with transcriptional dynamics in the following way:
$\mathcal{L}_{\mathrm{corr}}=-\left[\lambda_{u}\dfrac{\boldsymbol{\hat{v}}_{c}\cdot\boldsymbol{u}_{c}}{\|\boldsymbol{\hat{v}}_{c}\|\|\boldsymbol{u}_{c}\|}+\lambda_{s}\dfrac{\boldsymbol{\hat{v}}_{c}\cdot\boldsymbol{s}_{c}}{\|\boldsymbol{\hat{v}}_{c}\|\|\boldsymbol{s}_{c}\|}\right],$
@@
CellDancer Li et al. (2024) similarly parameterizes $\alpha_{g,c},\beta_{g,c},\gamma_{g,c}$ with neural networks, and estimates velocity as
$\boldsymbol{\hat{v}}_{c}=(\Delta\boldsymbol{u}_{c},\Delta\boldsymbol{s}_{c}),\quad\text{where }\Delta\boldsymbol{u}_{c}=\boldsymbol{\alpha}_{c}-\boldsymbol{\beta}_{c}\odot\boldsymbol{u}_{c},\ \Delta\boldsymbol{s}_{c}=\boldsymbol{\beta}_{c}-\boldsymbol{\gamma}_{c}\odot\boldsymbol{s}_{c}.$
Its loss maximizes velocity alignment with neighbors, as follows:
$\mathcal{L}=\sum_{c}\left[1-\max_{j\in\mathcal{N}(c)}\cos\left(\boldsymbol{\hat{v}}_{c},\boldsymbol{v}_{j}\right)\right],\quad\boldsymbol{v}_{j}=(\boldsymbol{u}_{j}-\boldsymbol{u}_{c},\boldsymbol{s}_{j}-\boldsymbol{s}_{c}).$
Vector Field Reconstruction Based on RNA Velocity After estimating the model parameters and obtaining the RNA velocity for each cell, an important downstream task is to predict the fate of cells based on the estimated velocity, e.g., the continuous differentiation trajectory of cells. To construct a continuous vector field, the RNA velocity of each cell could be treated as the value of a vector field at discrete points. Then, the vector field is reconstructed by formulating a regression problem. Subsequently, quantities derived from vector field analysis, such as equilibrium points, streamlines, gradients, divergence, and curl, are used to study cell fate in continuous dynamics setup Qiu et al. (2022).
@@
Denote the spliced RNA counts and RNA velocities at various data points $\{\boldsymbol{x}_{i},\boldsymbol{v}_{i}\}_{i=1}^{n}$. The task is to find a continuous vector field function $f^{\star}$ that minimizes the regression loss $\mathcal{L}=\sum_{i}p_{i}\|\boldsymbol{v}_{i}-\boldsymbol{f}(\boldsymbol{x}_{i})\|^{2}$ where $p_{i}$ denotes the weights of each data point.
Dynamo Qiu et al. (2022) approximates the unknown vector-valued function in a sparse reproducing kernel Hilbert space (RKHS). For a vector-valued function $\boldsymbol{f}\in\mathcal{H}$ in RKHS, it can be represented as a sum of Gaussian kernels as follows:
$\boldsymbol{f}(\boldsymbol{x})=\sum_{i=1}^{m}\Gamma(\boldsymbol{x},\boldsymbol{\tilde{x}}_{i})\boldsymbol{c}_{i},\quad\Gamma(\boldsymbol{x},\boldsymbol{\tilde{x}})=\exp(-w\|\boldsymbol{x}-\boldsymbol{\tilde{x}}\|^{2}),$
where $\boldsymbol{\tilde{x}}_{i}$ are called control points. Additionally, the norm of $\boldsymbol{f}$ in $\mathcal{H}$ can be computed as $\|\boldsymbol{f}\|^{2}=\sum_{i,j=1}^{m}\boldsymbol{c}_{i}^{T}\Gamma(\boldsymbol{\tilde{x}}_{i},\boldsymbol{\tilde{x}}_{j})\boldsymbol{c}_{j}^{T}.$ The loss function for the vector field estimation problem includes a regularization term based on the norm of $\boldsymbol{f}$ such that $\mathcal{L}_{\lambda}=\sum_{i}p_{i}\|\boldsymbol{v}_{i}-\boldsymbol{f}(\boldsymbol{x}_{i})\|^{2}+\frac{\lambda}{2}\|\boldsymbol{f}\|^{2},$ where $\lambda$ is the regularization coefficient.
Another popular fitting strategy to reconstruct the continuous vector field is to use the neural network, where a VAE-based deep learning method was proposed in Chen et al. (2022).
@@
Based on the estimated continuous vector field, Dynamo Qiu et al. (2022) proposed several analyses to reveal the differential geometry of the RNA velocity. First, the Jacobian matrix is essential for analyzing the stability of equilibrium points in a dynamical system and for studying gene–gene interactions. In RKHS context, the Jacobian matrix can be analytically computed as
$\boldsymbol{J}=\frac{\partial\boldsymbol{f}(\boldsymbol{x})}{\partial\boldsymbol{x}}=-2w\sum_{i=1}^{m}\Gamma(\boldsymbol{x},\boldsymbol{\tilde{x}}_{i})\boldsymbol{c}_{i}(\boldsymbol{x}-\boldsymbol{\tilde{x}}_{i})^{T}$
where $\Gamma(\boldsymbol{x},\boldsymbol{\tilde{x}}_{i})$ is the Gaussian kernel. Since the $(i,j)$-th entry of the Jacobian matrix $J_{ij}$ represents the effect of unspliced RNA levels of gene $j$ on the RNA velocity of gene $i$, the Jacobian matrix can be used to analyze the strength of gene–gene interactions. By averaging the Jacobian matrix across all data points, an average Jacobian matrix $\langle J\rangle$ can be obtained. By sorting the elements in each row of $\langle J\rangle$, the top regulators for each effector can be identified. Conversely, by sorting the elements in each column of $\langle J\rangle$, the top effectors for each regulator can be identified. The Jacobian matrix can also be used to compute the effect of perturbations. If the system state changes by $\Delta\boldsymbol{x}$ at $\boldsymbol{x}$, the resulting change in the vector field is $\Delta\boldsymbol{f}=\boldsymbol{J}\Delta\boldsymbol{x}.$
@@
    $\nabla\cdot\boldsymbol{f}=\mathrm{Tr}(\boldsymbol{J}),$
@@
    $\boldsymbol{a}=\frac{\mathrm{d}\boldsymbol{v}}{\mathrm{d}t}=\boldsymbol{J}\boldsymbol{v}.$
@@
    $\boldsymbol{\kappa}=\frac{1}{\|\boldsymbol{v}\|}\frac{\mathrm{d}}{\mathrm{d}t}\frac{\boldsymbol{v}}{\|\boldsymbol{v}\|}=\frac{(\boldsymbol{v}^{T}\boldsymbol{v})\boldsymbol{J}\boldsymbol{v}-(\boldsymbol{v}^{T}\boldsymbol{J}\boldsymbol{v})\boldsymbol{v}}{\|\boldsymbol{v}\|^{4}}.$
@@
Based on the learned vector field, continuous cell trajectories could also be constructed in Dynamo Qiu et al. (2022) using the concept of most probable path Zhou and Li (2016). Given the SDE (2), the action along any path $\Psi$ is defined as
$S_{T}[\psi]=\int_{0}^{T}L^{FW}\left(\psi,\dot{\psi}\right)dt,L^{FW}\left(\psi,\dot{\psi}\right)=\frac{1}{4}\left[\dot{\psi}(s)-b\left(\psi(s)\right)\right]^{t}D^{-1}\left(\psi(s)\right)\left[\dot{\psi}(s)-b\left(\psi(s)\right).\right]$
According to the Freidlin–Wentzell theorem Zhou and Li (2016), the path of least action is indeed the most probable path to make transitions between two attractors. In actual computation, Dynamo assumes the constant noise coefficient by taking $D=\frac{\sigma^{2}}{2}$ since only a continuous vector field is reconstructed.
@@
$R(\boldsymbol{x}_{s}\rightarrow\boldsymbol{x}_{t})\approx C\exp(-S^{\star}),$ (6)
@@ 3.2 Temporally Resolved Single-Cell RNA-Seq
Temporally resolved scRNA-seq provides us with a deeper understanding of the dynamics process in single cells. However, due to the destructive nature of scRNA-seq technology, we cannot track the trajectories of individual cells. Instead, we can only observe the changes in cellular distribution with time. Thus, reconstructing the trajectories of single cells from samples collected at discrete and sparse temporal points becomes crucial for understanding developmental processes and other dynamic biological processes and remains a challenging problem Schiebinger et al. (2019); Sha et al. (2024); Peng et al. (2024); Lavenant et al. (2024); Jiang et al. (2022); Jiang and Wan (2024); Bunne et al. (2024, 2023); Tong et al. (2023); Zhang et al. (2021); Maddu et al. (2024); Eyring et al. (2024); Zhang et al. (2024); Yeo et al. (2021). To overcome these challenges, many methods have been developed in recent years. From a dynamical perspective, these approaches can be broadly classified into two categories: those that model dynamics on discrete cell states and those that model dynamics in continuous spaces. In the following, we will introduce these methods separately from these two viewpoints. Figure 3 summarizes these approaches.
{adjustwidth}
[…]
Figure 3: Dynamic modeling of temporally-resolved single-cell transcriptomics.
@@ 3.2.1 Discrete Temporal Dynamics Modeling
Among the methods that model dynamics on discrete cell states, pioneering work includes Waddington OT Schiebinger et al. (2019) and Moscot Klein et al. (2025). These approaches employ static optimal transport as the main tools.
Static Optimal Transport To formulate this problem, consider $\boldsymbol{X}\in\mathbb{R}^{N\times G}$ and $\boldsymbol{Y}\in\mathbb{R}^{M\times G}$ represent two unpaired datasets of $N$ and $M$ cells observed at different time points ($t_{1},t_{2}$), respectively, in the $G$ dimensional gene expression space. Then, one can define two marginal distributions $\nu_{0}\in\mathcal{C}_{N}$, $\nu_{1}\in\mathcal{C}_{M}$ at $t_{1}$ and $t_{2}$ respectively on the probability simplex $\mathcal{C}_{N}=\{\boldsymbol{a}\in\mathbb{R}^{N}|\sum_{i=1}^{N}a_{i}=1,\boldsymbol{a}\geq 0\}.$ The goal of optimal transport is to find the optimal coupling $\boldsymbol{\pi}\in\mathbb{R}_{+}^{N\times M}$ that transports a distribution to another while minimizing the cost associated with the transportation. The feasible transport plan is defined as $\boldsymbol{\boldsymbol{\Pi}}(\nu_{0},\nu_{1})=\left\{\boldsymbol{\pi}\in\mathbb{R}^{N\times M}:\boldsymbol{\pi}\boldsymbol{1}_{M}=\nu_{0},\boldsymbol{\pi}^{T}\boldsymbol{1}_{N}=\nu_{1},\boldsymbol{\pi}\geq 0\right\}.$ So the static optimal transport problem is formally defined as
$\min_{\boldsymbol{\pi}\in\boldsymbol{\Pi}(\nu_{0},\nu_{1})}\langle\boldsymbol{\pi},\boldsymbol{c}\rangle=\sum_{i,j}c_{i,j}\pi_{i,j}.$ (7)
The cost matrix $\boldsymbol{c}\in\mathbb{R}_{+}^{n\times m}$ defines the transportation cost between each pair of points, where $c_{ij}:=c(\boldsymbol{x}_{i},\boldsymbol{y}_{j})$ quantifies the expense of transferring a unit mass from the source point $\boldsymbol{x}_{i}$ to the target point $\boldsymbol{y}_{j}$. By solving the static optimal transport problem, one can determine the transport matrix that couples the two distributions. This static optimal transport can be effectively addressed using the Python Optimal Transport (POT) library Flamary et al. (2021).
Entropic Optimal Transport Additionally, to enhance the efficiency of solving the optimal transport problem, a regularized optimal transport approach is often introduced. The discrete entropy of a coupling matrix is defined as
@@
$\min_{\boldsymbol{\pi}\in\boldsymbol{\Pi}(\boldsymbol{\nu_{0}},\boldsymbol{\nu_{1}})}\langle\boldsymbol{\pi},\boldsymbol{c}\rangle-\varepsilon H(\boldsymbol{\pi})$ (8)
Since the objective is an $\varepsilon$-strongly convex function, problem (8) has a unique optimal solution. Using the KKT conditions, the solution to (8) is unique and has the form Peyré et al. (2019)
@@
for two unknown variables $(\boldsymbol{a},\boldsymbol{b})\in\mathbb{R}_{+}^{N}\times\mathbb{R}_{+}^{M}$, where $K_{i,j}=e^{-\frac{c_{i,j}}{\varepsilon}}$.
Sinkhorn Algorithm From above, the optimal solution of problem (8) can be expressed in matrix form as $\boldsymbol{\pi}=\operatorname{diag}(\boldsymbol{a})\boldsymbol{K}\operatorname{diag}(\boldsymbol{b})$. Then it is necessary to satisfy the constraints $\boldsymbol{\Pi}(\nu_{0},\nu_{1})$, i.e.,
$\operatorname{diag}(\boldsymbol{a})\boldsymbol{K}\operatorname{diag}(\boldsymbol{b})\boldsymbol{1}_{M}=\nu_{0},\quad\operatorname{diag}(\boldsymbol{a})\boldsymbol{K}^{T}\operatorname{diag}(\boldsymbol{b})\boldsymbol{1}_{N}=\nu_{1}.$
Note that $\operatorname{diag}(\boldsymbol{b})\boldsymbol{1}_{M}$ is $\boldsymbol{b}$ and the multiplication of $\operatorname{diag}(\boldsymbol{a})$ times $\boldsymbol{K}\boldsymbol{b}$ is
$\boldsymbol{a}\odot(\boldsymbol{K}\boldsymbol{b})=\nu_{0},\boldsymbol{a}\odot\left(\boldsymbol{K}^{\mathrm{T}}\boldsymbol{b}\right)=\nu_{1},$
where $\odot$ represents the entry-wise multiplication. An intuitive way is to solve them iteratively, and these two updates yield the Sinkhorn algorithm,
$\boldsymbol{a}^{(\ell+1)}\stackrel{{\scriptstyle\text{ def. }}}{{=}}\frac{\nu_{0}}{\boldsymbol{K}\boldsymbol{b}^{(\ell)}},\quad\quad\boldsymbol{b}^{(\ell+1)}\stackrel{{\scriptstyle\text{ def. }}}{{=}}\frac{\nu_{1}}{\boldsymbol{K}^{{T}}\boldsymbol{a}^{(\ell+1)}}.$